OpenStudy (anonymous):

Write a system of two equations using x and y to represent the following problem. A chemist is mixing two solutions together. The chemist wants 100 gallons when complete. He is mixing a 24% solution with a 50% solution to get a 36% solution. How many gallons of each will he use?

OpenStudy (anonymous):

call the amount of 50% solution x and then the amount of 36% solution is 100-x

OpenStudy (anonymous):

oops let me start again. call the amount of 50% solution x and the amount of 24% solution is then 100-x

OpenStudy (anonymous):

so \[.5x+.24(100-x)=.36\times 100=36\]

OpenStudy (anonymous):

this is what you need to solve. \[.5x+24-.24x=36\] \[.26x+24=36\] \[.26x=12\] \[x=\frac{12}{.26}=46.15\] rounded

OpenStudy (anonymous):

but if i would read correctly it only says "wrote two equations" the first would be \[x+y=100\] and the second would be \[.5x+.24y=.6\]

OpenStudy (anonymous):

rather \[.5x+.24y=36\]

OpenStudy (anonymous):

Ah i see (:

OpenStudy (anonymous):

So then...

OpenStudy (anonymous):

it doesn't say to solve them does it?

OpenStudy (anonymous):

x+y = 100 .24x + .5y = .36 * 100

OpenStudy (anonymous):

Yea i do need to solve them. i can prolly do that on my own tho.... :P retaking assignments from the begining of the year >.< Dont remember a thing

OpenStudy (anonymous):

if you want to solve it a systems you have \[x+y=100\] \[.5x+.24y=36\]

OpenStudy (anonymous):

multiply the bottom equation by -2 and the x's will drop out

OpenStudy (anonymous):

\[x+y=100\] \[-x-.48y=-72\]

OpenStudy (anonymous):

dude, thats weird. let y = 100 - x and substitute that into the second equation,

OpenStudy (anonymous):

add them and the x's drop out to get one equation in y \[.52y=28\] \[y=\frac{28}{.52}=53.846\] rounded

OpenStudy (anonymous):

Naw he is doing elimination, thats substitution... i suck at substitution.. so go ahead and do elimination.. lol

OpenStudy (anonymous):

i did that first.

OpenStudy (anonymous):

Okay (:

OpenStudy (anonymous):

cantorset first line i used 100-x for y. but what is odd is that usually these problems are cooked up to give nice integer answers. this one is not.

OpenStudy (anonymous):

unless my arithmetic is wrong. i got two ugly decimal answer. it is 50%, 24% and 36% yes?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i trust your work , but first you want to multiply both sides to remove the decimal, if youre doing elimination

OpenStudy (anonymous):

for the second equation, multiply both sides by 100 first

OpenStudy (anonymous):

you try it and see if you don't get something ugly. maybe i made a mistake. really please try it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

still get \[x=\frac{12}{.26}=\frac{1200}{26}=46.153846...\]

OpenStudy (anonymous):

sat, bear with me, im tutoring 3 other people, sorry

OpenStudy (anonymous):

Oh yea! ^^ i think you said something about being a online tutor for money or something :P Continue though satellite (:

OpenStudy (anonymous):

and they COMPLAIN, lol

OpenStudy (anonymous):

yes real coin

OpenStudy (anonymous):

They people your tutoring? Haha

OpenStudy (anonymous):

You must be making bank.... lmfao.. kidding (:

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

not really, im working like a dog

OpenStudy (anonymous):

Wow... i sure cant wait to get done with high school..... :P lol

OpenStudy (anonymous):

So i can go on to work like a dog... and make no money... sounds wonderful :P lol

OpenStudy (anonymous):

still get x=46.15 rounded making y = 53.85 rounded

OpenStudy (anonymous):

Okay (: Thanks!

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

why are you doing this and also \[b^0\] at the same time? odd

OpenStudy (anonymous):

Haha i am retaking stuff i learned at the begining of the year :P I seriously barely remember anything. lol

OpenStudy (anonymous):

good luck. it will be over soon enuf

OpenStudy (anonymous):

Ya i sure hope so :P At least i got this cite tho! ^^

OpenStudy (anonymous):

what is b^0 ?

OpenStudy (anonymous):

Idk how to explain it... ^_T just know im doing it :P I bet satellite can tho (:

OpenStudy (anonymous):

wait, he said youre "taking" b^0

OpenStudy (anonymous):

Haha no its just something him and i think its polpak showed me how to do.

OpenStudy (anonymous):

Lilly allen sucks,